# Optimizing the potential of KM3NeT in detecting core-collapse supernovae

### v1

@@ B.2.2 Modeling atmospheric muons
Muon energies, entry points, and momenta are simulated using MUPAGE v1.5.1 [mupage1, mupage2]. The muon entry points are generated on the surface of a cylinder surrounding the detector. The radius and height of this cylinder are chosen such that muons passing up to $140\text{\,}\mathrm{m}$ ($280\text{\,}\mathrm{m}$) away from ORCA (ARCA) are simulated, with contributions from other muons amounting to less than $0.5\%$ of the single-DOM event rate for multiplicities $6$ and above. The muon propagation and detection is simulated using the JSirene software [jsirene1], which models the Cherenkov light emission induced by muons using tabulated probability density functions instead of tracking each photon individually. The high computational efficiency of JSirene makes it possible to model muon backgrounds for multiple ARCA and ORCA geometries with minimal computational cost.

### v2

@@ B.2.2 Modeling atmospheric muons
Muon energies, entry points, and momenta are simulated using MUPAGE v1.5.1 [11, 14]. The muon entry points are generated on the surface of a cylinder surrounding the detector. The radius and height of this cylinder are chosen such that muons passing up to $140\text{\,}\mathrm{m}$ ($280\text{\,}\mathrm{m}$) away from ORCA (ARCA) are simulated, with contributions from other muons amounting to less than $0.5\%$ of the single-DOM event rate for multiplicities $6$ and above. The muon propagation and detection is simulated using the JSirene software [18], which models the Cherenkov light emission induced by muons using tabulated probability density functions instead of tracking each photon individually. The high computational efficiency of JSirene makes it possible to model muon backgrounds for multiple ARCA and ORCA geometries with minimal computational cost.